What is the percent composition of the elements CuBr2
1
Expert's answer
2012-12-19T10:42:28-0500
Mw of CuBr2 = 63.546 + 79.904*2 = 223.354

% of Cu = 63.546/223.354*100% = 28.451%
% of Br = 79.904*2/223.354* 100% = 71.549%